OpenStudy (anonymous):

2.Two bicycles depart from Miami Beach going in opposite directions. The first bicycle is traveling at 10 miles per hour. The second bicycle travels at 5 miles per hour. How long does it take until the bikes are 45 miles apart?

OpenStudy (anonymous):

D=RT 45=(10+5)T 45=15T T=445/15 T=3 HOURS ANS.

OpenStudy (anonymous):

ok you are home schooled and go to connex or k-12... i di the same thing all the time

OpenStudy (anonymous):

yea at the moment i am. thank you